# Image-to-Video Transfer Learning based on Image-Language Foundation Models: A Comprehensive Survey

### v1

@@ IV-C Temporal Video Grounding
Beyond the post-network paradigm, other transfer strategies have been explored. VDI [60] proposes a language-side adaptation paradigm for video-language models. While keeping the CLIP visual encoder frozen, it adapts the text encoder by infusing it with video-derived supervisory signals via a side-tuning approach. R2-Tuning [59] introduced a Side-Tuning method in the form of a reversed recurrent tuning block, which performs query-modulated spatial pooling and coarse-to-fine temporal refinement. To incorporate explicit relational reasoning, PaTF[61] encodes external cues by introducing a parallel transformer that processes structured scene graph representations.
@@
More recently, the advent of Multimodal Large Language Models has introduced a new research direction focused on enhancing their temporal awareness. LLaViLo[108] was the first to utilize a video semantic modeling adapter to inject a joint feature representation into an LLM. To incorporate explicit time information as external cues, Chrono [96] introduced a sequence-to-sequence approach that integrates video frames with their corresponding timestamp cues. Similarly, VTG-LLM [100] integrated timestamp knowledge through sequence-time embeddings, absolute-time tokens, and slot-based token compression. ReVisionLLM [109] incorporated a recursive processing architecture with hierarchical adapters to progressively refine temporal segments in hour-long videos. In a training-free manner, TFVTG [48] employed LLMs to reason about sub-event relationships while using separate VLMs to localize each sub-event. NumPro [95] employs a fine-tuning strategy augmented with explicit visual cues. It overlays unique numerical identifiers onto each video frame, thereby translating the temporal grounding task into a visual OCR problem that can be directly resolved by a Video LLM.
IV-D Spatio-Temporal Video Grounding
Spatio-Temporal Video Grounding aims to predict a spatio-temporal tube for the event depicted in given language query from an untrimmed video. HCSTVG-v1, HCSTVG-v2, VidSTG-Declarative, and VidSTG-Interrogative are the most popular benchmarking datasets used in the community. The evaluation metrics for this task include: (1) m_vIoU, vIoU@0.3, and vIoU@0.5, which measure the spatiotemporal localization accuracy; and (2) m_tIoU, which solely evaluates the temporal localization precision.
@@ IV-D Spatio-Temporal Video Grounding
Earlier methods [157, 158, 159] use unimodal detectors like Faster R-CNN [160] to detect objects in each video frame and then ground the temporal locations based on the detection features. STVGBert [68] is the first work to introduce and fully finetune a pre-trained ViLBERT [161] model and develop a novel spatio-temporal composition/decomposition attention module, thereby enabling joint spatio-temporal modeling. Recent studies [70, 71, 89, 162, 90, 92, 91] have developed frameworks based on the powerful pre-trained MDETR model [163], incorporating additional spatio-temporal interaction mechanisms. To achieve better performance, some approaches [70, 71] additionally introduce a temporal module to capture the cross-frame interaction and fine-tune all the parameters involved in the model. Specifically, TubeDETR [70] introduces a fast visual-only branch in the encoder and a time-aligned cross-attention mechanism in the decoder, enabling object queries to achieve joint spatio-temporal modeling. STCAT [71] designs a video-level multimodal template in the decoder to maintain temporal consistency across spatial predictions.
@@
In weakly-supervised STVG, the recent SOTA methods have increasingly adopted post-network tuning strategies. The statical vision-language alignment knowledge learned by pre-trained grounding models (like Grounding-DINO [142]) is explicitly exploited in the community, alleviating the requirement of manual spatiotemporal annotations in the training stage. For example, CoSPaL [49] and STPro [50] leverage pretrained Grounding-DINO to generate frame-wise spatial predictions, which are then concatenated to form a set of candidate tubelets with annotations. They then develop a curriculum learning strategy to capture spatio-temporal correspondences between sentence and tubelets, filtering out non-relevant proposals.
@@ V Coarse-grained Video Understanding
Compared to fine-grained video-text understanding, coarse-grained tasks do not require precise temporal grounding results. They place greater emphasis on the holistic comprehension of events depicted in both text and video, such as determining whether a text-described event occurs in a video (Video Question Answering) or generating textual descriptions of video content (Video Captioning). In this section, we review methods for coarse-grained video-text understanding tasks from the perspective of transferring Image-Language Foundation Models to video domains.
@@ V-A Video-Text Retrieval
Video–Text Retrieval (VTR) aims to retrieve the most relevant video from a large corpus given a textual query, or vice versa. It serves as a fundamental task in video understanding and often acts as a precursor to more fine-grained video–language tasks. Since the introduction of the large-scale image–text pretrained model CLIP [1], a series of subsequent studies have focused on transferring its knowledge to video–text retrieval, achieving substantial performance gains over earlier methods. Early approaches in this domain typically adopt a full fine-tuning paradigm, where model parameters are initialized from pretrained weights and subsequently optimized for downstream tasks. that explored transferring CLIP’s image representations to the video domain for video–text retrieval. It employs a 2D or 3D linear projection layer on flattened patches and introduces an additional learnable token to extract frame-level embeddings, following a pipeline similar to that of the vanilla ViT [165]. In addition, a preliminary video-text post-pretraining approach was explored in CLIP4Clip to further adapt CLIP representations to the video domain. CLIP2Video [63] extends image–text pretrained models to video–text retrieval by introducing a temporal difference block and a temporal alignment block. The temporal difference block enhances motion-related cues by modeling differences between adjacent frames, while the temporal alignment block aggregates text and frame embeddings through learnable center embeddings in the joint space. Another method [80] improves the full fine-tuning performance of CLIP for VTR by developing a weighted token-wise interaction mechanism and a channel de-correlation regularization loss. CLIP-Hitchhiker [78] observed that most single-modal temporal modeling and aggregation strategies under-perform a simple average temporal pooling operation. To address this, it introduced a query-guided temporal aggregation baseline, which achieved performances superior to state-of-the-art methods. CenterCLIP [81] proposes a more efficient full fine-tuning paradigm to transfer CLIP’s knowledge, where token clustering is conducted within each video segment and only center tokens are retained and further processed to compute the video-level embedding. TS2-Net [115] combines a temporal token shifting operation and a spatial token selection module to enhance the spatial-temporal video representation learning for better retrieval. CLIP2TV [35] constructs soft labels from the intra-modality similarity matrices to guide the learning of a fusion-based video-text matching module. X-CLIP [124] transfers the knowledge of pretrained CLIP models to end-to-end video–text retrieval through multi-grained contrastive learning across video–sentence and frame–word representations. In addition, an Attention over Similarity Matrix (AOSM) mechanism is employed to integrate similarity scores across multiple cross-modal levels. CLIP-ViP [126] introduces video proxy tokens as intermediaries for local–global information exchange, and further conducts video–text post-pretraining followed by downstream fine-tuning. To mitigate the injection of irrelevant visual information during video representation learning, X-pool [101] introduces a text-conditioned frame embedding pooling module, enabling more effective transfer of CLIP for video–text retrieval. VOP [125] introduces a cooperative text–video prompt-tuning strategy to mitigate catastrophic forgetting of pretrained vision–language knowledge while substantially reducing the number of trainable parameters. Considering the mutual dependency between contrastive dual encoders and the spatio-temporal characteristics of video data, VOP inserts learnable prompt tokens at the inputs of each text and vision encoder block, with positional, contextual, and functional aspects of video prompts explicitly modeled to enable effective transfer. CLIPPING [38] proposes an efficient knowledge distillation framework with a novel Student-As-Base (SAB) scheme, which regards the student’s intermediate features as the bases of the teacher’s feature space to absorb the pretrained knowledge. PromptSwitch [166] inserts a cube tensor as learnable prompts into the CLIP image encoder, where the prompt cube can be flexibly transposed to assist in the peer-to-peer communication between each frame pairs in the video. RAP [167] designs a low-rank modulation operation in the adapter module to inject the temporal sparsity and correlation characteristics for efficient image-to-video transfer. MV-Adapter [168] equips the adapter module with two novel components of temporal adaptation and cross modality typing that can effectively tackle the task of parameter-efficient video-text retrieval. TeachCLIP [39] introduces a multi-grained knowledge distillation framework that leverages a frozen CLIP model as the teacher. It jointly performs fine-grained frame-level distillation and coarse-grained video-level distillation, leading to strong performance in video–text retrieval. DiscoVLA [119] identifies three key image-to-video transfer gaps including vision, language and alignment and then introduces the IVFusion and PImgAlign to accordingly address the three transfer gaps for parameter-efficient video-text retrieval.
@@ V-B Video Action Recognition
Video action recognition is a classical task in video understanding, which aims to recognize the category of action being performed in trimmed video. Over the years, there have been plentiful research works making efforts to transfer the powerful pretrained image-text foundation models for tackling video action recognition. ActionCLIP [121] is a pioneering work that adopts a “pretrain, prompt, and fine-tune” paradigm for transferring the knowledge of CLIP, which formulates action recognition into a multimodal learning problem by extracting text representations of class labels for video-text matching. X-CLIP [116] introduces a new framework to expand existing large-scale image-text pretrained models for general video recognition, where a cross-frame communication attention module and a video-specific textual prompting strategy are incorporated for better transfer performance. EVL [65] proposes to utilize a sideway transformer decoder to progressively aggregate multiple intermediate features from a frozen pretrained backbone to a learnable query token that is regarded as the global feature for predicting action classes. ST-Adapter [111] presents an efficient spatio-temporal adapter which is instantiated as a 3D convolution module inserted at the front end of each pretrained block in cascade. Adaptformer [112] designs another kind of adapter module which is inserted as a parallel learnable branch for merging into the original MLP’s output. AIM [113] jointly incorporates cascaded adapters into the spatial and temporal attention layers and adds an adapter in parallel to the MLP layer to achieve efficient transfer learning. BIKE [106] explores a bidirectional cross-modal knowledge transfer approach, where the attributes-category association and video concept spotting mechanisms are introduced to complement the unidirectional video-to-text matching and enhance the temporal awareness of video representations. ViFi-CLIP [79] shows that simple full fine-tuning of pretrained CLIP models with late temporal pooling generalizes well to supervised settings and a vision-language prompting scheme works well in low-data regimes. Text4Vis [36] demonstrates that a new paradigm with frozen classifiers initialized from pretrained textual embeddings can achieve more efficient transfer learning as compared to previous contrastive-based paradigms. DiST [66] disentangles spatial and temporal modeling to the pretrained backbone and a learnable network, and further utilizes a side branch to integrate multiple intermediate features from the spatial and temporal streams. UniFormerV2 [7] designs a sophisticated transfer method by additionally learning local temporal modeling and global spatio-temporal modeling on top of the pretrained global spatial modeling and a multi-stage fusion block is utilized to aggregate video tokens from different semantic levels. DUALPATH [114] explicitly separates the image-to-video adaptation into a spatial path with parallel adapters and a temporal path with serial adapters followed by a late path fusion for action classification. In addition, a new manner to imitate spatial modeling to achieve efficient temporal modeling via token interactions in a grid-like frameset is also proposed. MoMa [102] proposes a new temporal modeling method called sequence modulation operation, which is achieved by efficient trainable mamba layers to inject fine-grained spatio-temporal information into intermediate features. D2ST [117] develops a dual-pathway adapter that enables disentangled adaptation of spatial and temporal features, while maintaining full compatibility with pretrained single-stream image-text models.
@@ V-C Video Captioning
Video captioning aims to generate natural language descriptions of video content. It has evolved from early template-based methods [169, 170, 171] to the adaptation of large pre-trained foundation models. Many works adopt a post-network tuning strategy to enhance features from frozen image encoders. Clip4Caption [56] introduces a post-network for temporal segment sampling and utilizes contrastive pretraining to learn text-aligned video embeddings from CLIP features. EvCap [105] also employs a post-network approach to fuse global visual contexts with specific action and object cues extracted by a pretrained CLIP model. mPLUG-2 [87] introduces a partial pretraining paradigm where a dual-vision encoder decouples spatial and temporal modeling by augmenting shared spatial self-attention layers with a novel local temporal modeling module for video inputs. MAMS [67] introduces a model-agnostic Side-Tuning framework that utilizes a module selector to adaptively choose between captioning modules of different sizes and incorporates an adaptive attention masking scheme to focus on important visual tokens. Similarly, CM2 [104] enhances CLIP features using multi-scale temporal convolutions and a cross-modal memory module that retrieves relevant textual semantics as external cues. In contrast, SWINBERT [69] proposes an end-to-end fine-tunable architecture, employing a Video Swin Transformer to extract dense spatio-temporal tokens which are then processed by a multimodal transformer with sparse attention for effective long-range modeling. More recent approaches focus on improving temporal precision. Vid2Seq [103] presents a unified framework for joint event captioning and localization by augmenting the language model’s vocabulary with special time tokens that serve as explicit temporal models. VTG-LLM [100] enhances pretrained encoders by injecting temporal cues like sequential and absolute time embeddings and employs a slot-based token compression mechanism for the efficient handling of long video sequences. HierarQ [58] leverages a trainable hierarchical Q-Former as post network, equipped with a two-stream feature modulator and dedicated memory banks, to compress frozen video features into a compact representation for a LoRA-adapted LLM.

### v2

@@ IV-B Temporal Video Grounding
Beyond the post-network paradigm, other transfer strategies have also been explored. VDI [125] introduces a language-side adaptation paradigm in which the CLIP visual encoder remains frozen, while the text encoder is adapted by infusing video-derived supervisory signals through a side-tuning approach. R2-Tuning [118] implements side-tuning using a reversed recurrent tuning block to explicitly perform query-modulated spatial pooling and coarse-to-fine temporal refinement. To incorporate explicit relational reasoning, PaTF [134] encodes external cues by introducing a parallel transformer that processes structured scene graph representations.
@@
More recently, Multimodal Large Language Models are introduced to enhance temporal reasoning. LLaViLo [128] utilizes a video semantic modeling adapter to inject a joint feature representation into an LLM. To incorporate explicit time information as external cues, Chrono [131] introduced a sequence-to-sequence approach that integrates video frames with their corresponding timestamp cues. Similarly, VTG-LLM [58] integrated timestamp knowledge through sequence-time embeddings, absolute-time tokens, and slot-based token compression. ReVisionLLM [59] incorporated a recursive processing architecture with hierarchical adapters to progressively refine temporal segments in hour-long videos. In a training-free manner, TFVTG [213] employed LLMs to reason about sub-event relationships while using separate VLMs to localize each sub-event. NumPro [185] employs a fine-tuning with extra models strategy to achieve video-language understanding. It overlays unique numerical identifiers onto each video frame, thereby translating the temporal grounding task into a visual OCR problem that can be directly resolved by a Video LLM.
IV-C Spatio-Temporal Video Grounding
@@
STVG aims to predict a spatio-temporal tube for the event depicted in given language query from an untrimmed video. Earlier methods [210, 209, 162] use unimodal detectors like Faster R-CNN [149] to detect objects in each video frame and then ground the temporal locations based on the detection features. STVGBert [155] is the first work to introduce and fully finetune a pre-trained ViLBERT [123] model and develop a novel spatio-temporal composition/decomposition attention module, thereby enabling joint spatio-temporal modeling. Recent studies [195, 71, 110, 175, 105, 56, 55] have developed frameworks based on the powerful pre-trained MDETR model [74], incorporating additional spatio-temporal interaction mechanisms. To achieve better performance, some approaches [195, 71] additionally introduce a temporal module to capture the cross-frame interaction and fine-tune all the parameters involved in the model. Specifically, TubeDETR [195] introduces a fast visual-only branch in the encoder and a time-aligned cross-attention mechanism in the decoder, enabling object queries to achieve joint spatio-temporal modeling. STCAT [71] designs a video-level multimodal template in the decoder to maintain temporal consistency across spatial predictions.
@@
IV-D Video Segmentation
Open Vocabulary Video Instance Segmentation (OV-VIS) extends traditional VIS by enabling the segmentation and tracking of of unseen object categories, leveraging vision-language foundation models to generalize beyond a predefined label set. Current research mainly leverages CLIP’s open-vocabulary capabilities for this task, designing additional network modules after the CLIP encoder to enable image-to-video transfer. OV2Seg [169] is the first to integrate CLIP into a post-network fine-tuning strategy for OVVIS. It adopts a DETR-like architecture with object-centric queries for mask prediction and cross-frame matching. The query embeddings, which combine information from current and previous frames, are compared with CLIP-extracted text features for open-vocabulary classification. InstFormer [57] integrates instance-guided attention into CLIP’s visual encoder, using segmentation cues to aggregate text-aligned features, which are then fed into the tracking module for instance matching across frames. BriVIS [26] uses a frozen video instance segmentor to generate frame-level queries and integrates multi-scale CLIP visual features for temporal modeling. It introduces a Brownian bridge to constrain instance features, using a contrastive objective to align the bridge center with the correct class text embedding while distancing it from irrelevant classes. OVFormer [39] constructs video-level instance queries and aligns them with frozen CLIP features using cross-attention for unified segmentation and tracking across frames. OV-DVIS++ [206] propose a address the video segmentation problem via a segment-tracking-refinement procedure in a local to global manner. CLIP-VIS [217] directly leverages CLIP’s image features to generate masks by combining predicted object scores and mask IoU. To alleviate the effect of historical noise, it first applies Hungarian algorithm to identify the $K$ most relevant frames from memory, and then performs instance tracking.
@@ V Coarse-grained Video Understanding
Compared to fine-grained video-text understanding, coarse-grained tasks do not require precise temporal grounding results. They place greater emphasis on the holistic comprehension of events depicted in both text and video, such as determining whether a text-described event occurs in a video (Video QA) or generating textual descriptions of video content (Video Captioning). In this section, we review methods for coarse-grained video-text understanding tasks.
@@ V-A Video-Text Retrieval
Video–Text Retrieval (VTR) aims to retrieve the most relevant video from a large corpus given a textual query, or vice versa. Since the introduction of the large-scale image–text pretrained model CLIP [146], a series of subsequent studies have focused on transferring its knowledge to video–text retrieval, achieving substantial performance gains over earlier methods. Early approaches in this domain typically adopt a full fine-tuning paradigm, where model parameters are initialized from pretrained weights and subsequently optimized for downstream tasks. CLIP4Clip [126] is the first method that explored transferring CLIP’s image representations to the video domain for video–text retrieval. It employs a 2D or 3D linear projection layer on flattened patches and introduces an additional learnable token to extract frame-level embeddings, following a pipeline similar to that of the vanilla ViT [35]. In addition, a preliminary video-text post-pretraining approach was explored in CLIP4Clip [126] to further adapt CLIP representations to the video domain. CLIP2Video [38] extends image–text pretrained models to video–text retrieval by introducing a temporal difference block and a temporal alignment block. The temporal difference block enhances motion-related cues by modeling differences between adjacent frames, while the temporal alignment block aggregates text and frame embeddings through learnable center embeddings in the joint space. Another method [173] improves the full fine-tuning performance by developing a weighted token-wise interaction mechanism and a channel de-correlation regularization loss.
@@
CLIP-Hitchhiker [5] observed that most single-modal temporal modeling and aggregation strategies under-perform a simple average temporal pooling operation. To address this, it introduced a query-guided temporal aggregation baseline, which achieved performances superior to SOTA methods. CenterCLIP [211] proposes a more efficient full fine-tuning paradigm to transfer CLIP’s knowledge, where token clustering is conducted within each video segment and only center tokens are retained and further processed to compute the video-level embedding. TS2-Net [120] combines a temporal token shifting operation and a spatial token selection module to enhance the spatial-temporal video representation learning for better retrieval. CLIP2TV [47] constructs soft labels from the intra-modality similarity matrices to guide the learning of a fusion-based video-text matching module. X-CLIP [129] transfers the knowledge of pretrained CLIP models to end-to-end video–text retrieval through multi-grained contrastive learning across video–sentence and frame–word representations. In addition, an Attention over Similarity Matrix (AOSM) mechanism is employed to integrate similarity scores across multiple cross-modal levels. CLIP-ViP [193] introduces video proxy tokens as intermediaries for local–global information exchange, and further conducts video–text post-pretraining followed by downstream fine-tuning. To mitigate the injection of irrelevant visual information during video representation learning, X-pool [53] introduces a text-conditioned frame embedding pooling module, enabling more effective transfer of CLIP. VOP [65] introduces a cooperative text–video prompt-tuning strategy to mitigate catastrophic forgetting of pretrained vision–language knowledge while substantially reducing the number of trainable parameters. Considering the mutual dependency between contrastive dual encoders and the spatio-temporal characteristics of video data, VOP inserts learnable prompt tokens at the inputs of each text and vision encoder block, with positional, contextual, and functional aspects of video prompts explicitly modeled to enable effective transfer.
CLIPPING [141] proposes an efficient knowledge distillation framework with a novel Student-As-Base (SAB) scheme, which regards the student’s intermediate features as the bases of the teacher’s feature space to absorb the pretrained knowledge. PromptSwitch [31] inserts a cube tensor as learnable prompts into the CLIP image encoder, where the prompt cube can be flexibly transposed to assist in the peer-to-peer communication between each frame pairs in the video. RAP [10] designs a low-rank modulation operation in the adapter module to inject the temporal sparsity and correlation characteristics for efficient image-to-video transfer. MV-Adapter [70] equips the adapter module with two novel components of temporal adaptation and cross-modality typing that can effectively tackle the task of parameter-efficient video-text retrieval. TeachCLIP [163] introduces a multi-grained knowledge distillation framework to perform fine-grained frame-level distillation and coarse-grained video-level distillation, leading to strong performance in video–text retrieval. DiscoVLA [152] introduces parameter-efficient fusion and distillation modules to simultaneously mitigate the vision, language, and alignment gaps in existing image-to-video transfer learning paradigms.
@@ V-B Video Action Recognition
Video action recognition is a classical task in video understanding, which aims to recognize the category of action being performed in a trimmed video. Over the years, there have been plentiful research works making efforts to transfer the powerful pretrained ILFMs for tackling video action recognition. ActionCLIP [171] is a pioneering work that adopts a “pretrain, prompt, and fine-tune” paradigm for transferring the knowledge of CLIP, which formulates action recognition into a multimodal learning problem by extracting text representations of class labels for video-text matching. X-CLIP [135] introduces a new framework to expand existing large-scale image-text pretrained models for general video recognition, where a cross-frame communication attention module and a video-specific textual prompting strategy are incorporated. EVL [111] proposes to utilize a sideway transformer decoder to progressively aggregate multiple intermediate features from a frozen pretrained backbone to a learnable query token that is regarded as the global feature for predicting action classes. ST-Adapter [137] presents an efficient spatio-temporal adapter which is instantiated as a 3D convolution module inserted at the front end of each pretrained block in cascade.
Adaptformer [22] designs another kind of adapter module which is inserted as a parallel learnable branch for merging into the original MLP’s output. AIM [197] jointly incorporates cascaded adapters into the spatial and temporal attention layers and adds an adapter in parallel to the MLP layer to achieve efficient transfer learning. BIKE [184] explores a bidirectional cross-modal knowledge transfer approach, where the attributes-category association and video concept spotting mechanisms are introduced to complement the unidirectional video-to-text matching and enhance the temporal awareness of video representations. ViFi-CLIP [147] shows that simple full fine-tuning of pretrained CLIP models with late temporal pooling generalizes well to supervised settings and a vision-language prompting scheme works well in low-data regimes. Text4Vis [183] demonstrates that a new paradigm with frozen classifiers initialized from pretrained textual embeddings can achieve more efficient transfer learning as compared to previous contrastive-based paradigms.
DiST [145] disentangles spatial and temporal modeling to the pretrained backbone and a learnable network, and further utilizes a side branch to integrate multiple intermediate features from the spatial and temporal streams. UniFormerV2 [92] designs a sophisticated transfer method by additionally learning local temporal modeling and global spatio-temporal modeling on top of the pretrained global spatial modeling and a multi-stage fusion block is utilized to aggregate video tokens from different semantic levels. DUALPATH [140] explicitly separates the image-to-video adaptation into a spatial path with parallel adapters and a temporal path with serial adapters followed by a late path fusion for action classification. In addition, a new manner to imitate spatial modeling to achieve efficient temporal modeling via token interactions in a grid-like frameset is also proposed. MoMa [198] proposes a new temporal modeling method called sequence modulation operation, which is achieved by efficient trainable mamba layers to inject fine-grained spatio-temporal information into intermediate features. D2ST [142] develops a dual-pathway adapter that enables disentangled adaptation of spatial and temporal features, while maintaining full compatibility with pretrained single-stream image-text models.
@@ V-C Video Captioning
Video captioning aims to generate natural language descriptions of video content. It has evolved from early template-based methods [168, 66, 78] to the adaptation of large pre-trained foundation models. Many works adopt a post-network tuning strategy to enhance features from frozen image encoders. Clip4Caption [161] introduces a post-network for temporal segment sampling and utilizes contrastive pretraining to learn text-aligned video embeddings from CLIP features. EvCap [116] also employs a post-network approach to fuse global visual contexts with specific action and object cues extracted by a pretrained CLIP model. mPLUG-2 [189] introduces a partial pretraining paradigm where a dual-vision encoder decouples spatial and temporal modeling by augmenting shared spatial self-attention layers with a novel local temporal modeling module for video inputs. MAMS [84] introduces a model-agnostic Side-Tuning framework that utilizes a module selector to adaptively choose between captioning modules of different sizes and incorporates an adaptive attention masking scheme to focus on important visual tokens.
Similarly, CM2 [77] enhances CLIP features using multi-scale temporal convolutions and a cross-modal memory module that retrieves relevant textual semantics as external cues. In contrast, SWINBERT [108] proposes an end-to-end fine-tunable architecture, employing a Video Swin Transformer to extract dense spatio-temporal tokens which are then processed by a multimodal transformer with sparse attention for effective long-range modeling. More recent approaches focus on improving temporal precision. Vid2Seq [196] presents a unified framework for joint event captioning and localization by augmenting the language model’s vocabulary with special time tokens that serve as an explicit temporal modeling component. VTG-LLM [58] enhances pretrained encoders by injecting temporal cues like sequential and absolute time embeddings and employs a slot-based token compression mechanism for the efficient handling of long video sequences. HierarQ [2] leverages a trainable hierarchical Q-Former as post network, equipped with a two-stream feature modulator and dedicated memory banks, to compress frozen video features into a compact representation for a LoRA-adapted LLM.
[…]